# On The Structure of EFX Orientations on Graphs

### v1

@@ 6. Discussion
In this paper, we studied strong EFX-orientability for graphical valuations [Christodoulou2023]. We showed a deep connection of this property to the chromatic number of the graph: every strongly EFX-orientable graph has a chromatic number of at most three, and a graph with a chromatic number of two or less is strongly EFX-orientable. This result is tight in the following sense where we demonstrate a 3-chromatic graph with and without this property.
We demonstrated that 0-1 strong EFX orientability and general strong EFX orientability were separate notions, and it remains open as to whether there is a graph that is EFX orientable for all additive valuations but not EFX orientable for a monotone function, or if additive strong EFX-orientability and strong EFX-orientability are equivalent. When given a graph with the valuation functions, determining whether an EFX-orientation exists is NP-hard [Christodoulou2023], but it is currently unknown if determining whether a graph is strongly EFX-orientable or not is NP-hard. Finally, it would be interesting to characterize graphs that allow EFX orientation, which is also Pareto optimal (PO).

### v2

@@ 6. Discussion
In this paper, we studied strong EFX-orientability for graphical valuations [Christodoulou2023]. We showed a deep connection of this property to the chromatic number of the graph: every strongly EFX-orientable graph has a chromatic number of at most three, and a graph with a chromatic number of two or less is strongly EFX-orientable. This result is tight in the following sense where we demonstrate a 3-chromatic graph with and without this property. The upper bound of 3 on the chromatic number demonstrates that apart from a niche class of graphs with nice structure, most graphs are not strongly EFX orientable. Therefore, it is reasonable to expect that when forming an EFX allocation in this graphical setting, not all edges are allocated to an incident vertex that values them.
We also demonstrated that 0-1 strong EFX orientability and general strong EFX orientability were separate notions, and it remains open as to whether there is a graph that is EFX orientable for all additive valuations but not EFX orientable for a monotone function, or if additive strong EFX-orientability and strong EFX-orientability are equivalent. When given a graph with the valuation functions, determining whether an EFX-orientation exists is NP-hard [Christodoulou2023], but it is currently unknown if determining whether a graph is strongly EFX-orientable or not is NP-hard. It would be interesting to characterize graphs which admit an EFX orientation that is also Pareto optimal (PO). Finally, we hope that our results will help with obtaining a complete characterization of strongly EFX-orientable graphs.
Acknowledgements
The first author would like to thank Elfarouk Harb and Daniel Schoepflin for a careful reading of this manuscript, and Pooja Kulkarni for thoughtful discussions.